Reports 1-1 of 1 Clear search Modify search
AdV-INJ (Input Mode Cleaner cavity)
derossi, spinicelli, gosselin - 17:53 Thursday 20 November 2025 (68203) Print this report
inj instabilities since the chiller swap

Since yesterday evening some instabilities of the INJ system have been observed (see as ex. plot 1: just before the IMC unlock the EOM corrections increased).

We measured the IMC TF, finding a very low UGF (around 75kHz, fig. 2).

The only modification on the system we made yesterday has been the swap of the NEOVAN chiller (#68190), which reduced the temperature of the neovan and slightly modified the coupling in the IMC (see IMC REFL decrease, plot 3). Not sure if it is the only cause of this change in the loop behaviour. Moreover, the temperature seems less stable (tomorrow we will decide whether to swap again the chiller, which is now working again thanks to fabio and cecilia).

By decreasing the attenuation of the IMC PDH loop from 21dB to 16dB (i.e. increase 5dB the gain) we recovered a more usual UGF and phase margin (fig.4)

Images attached to this report
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×